# An Overview of Systems-Theoretic Guarantees in Data-Driven Model Predictive Control

### v1

@@ 6 DISCUSSION
When applying data-driven control to real-world systems, especially in complex and safety-critical applications, it is desirable to provide rigorous systems-theoretic guarantees for the closed-loop operation. In this review, we discussed direct data-driven MPC methods based on the Fundamental Lemma which do provide such guarantees, e.g., on stability, robustness, and constraint satisfaction. The presented schemes can control unknown systems based only on input-output data in various scenarios including linear and nonlinear systems as well as noise-free and noisy data. As in the model-based case, closed-loop stability of data-driven MPC is not necessarily guaranteed but needs to be ensured via suitable terminal ingredients or a sufficiently long prediction horizon. When controlling linear systems based on noise-free data, techniques from model-based MPC can be borrowed, but the technical analysis needs to additionally cope with input-output cost functions. In the more realistic scenario of noisy data (bounded or stochastic), desirable guarantees can be provided under appropriate modifications of the MPC scheme, e.g., adding a regularization. For nonlinear systems, we discussed two alternatives based on either global or local linearity. The former approach exploits that systems from specific classes can be represented in linear coordinates using knowledge of suitable basis functions, allowing to formulate nonlinear versions of the Fundamental Lemma that can be used for MPC. Alternatively, nonlinear systems can be controlled by updating the data used for prediction via the Fundamental Lemma online, exploiting local linearity of (smooth) nonlinear dynamics. In guaranteeing closed-loop properties for the presented approaches, the key challenge is to combine established concepts from model-based MPC with unique challenges when using the Fundamental Lemma for prediction, in particular inaccuracies due to noise or nonlinearities.
@@
    Transferring open-loop results to closed-loop formulations: Several effects in data-driven MPC are by now well-understood in open-loop optimal control formulations, but have not been researched extensively w.r.t. closed-loop guarantees, e.g., regularization. On the other hand, when applying data-driven MPC, typically the closed-loop performance is the main object of interest. Thus, transferring open-loop to closed-loop results is another promising future research avenue.

### v2

@@ 6 DISCUSSION
When applying data-driven control to real-world systems, especially in complex and safety-critical applications, it is desirable to provide rigorous systems-theoretic guarantees for the closed-loop operation. In this review, we discussed direct data-driven MPC methods based on the Fundamental Lemma which do provide such guarantees, e.g., on stability, robustness, and constraint satisfaction. The presented schemes can control unknown systems based only on input-output data in various scenarios including linear and nonlinear systems as well as noise-free and noisy data. As in the model-based case, closed-loop stability of data-driven MPC is not necessarily guaranteed but needs to be ensured via suitable terminal ingredients or a sufficiently long prediction horizon. When controlling linear systems based on noise-free data, techniques from model-based MPC can be borrowed, but the technical analysis needs to additionally cope with input-output cost functions. In the more realistic scenario of noisy data (bounded or stochastic), desirable guarantees can be provided under appropriate modifications of the MPC scheme, e.g., adding a regularization. For nonlinear systems, we discussed two alternatives based on either global or local linearity. The former approach exploits that systems from specific classes can be represented in linear coordinates using suitable basis functions, leading to nonlinear versions of the Fundamental Lemma that can be used for MPC. Alternatively, nonlinear systems can be controlled by updating the data used for prediction via the Fundamental Lemma online, exploiting that (smooth) nonlinear dynamics are locally well-approximated by affine dynamics. In guaranteeing closed-loop properties for the presented approaches, the key challenge is to combine established concepts from model-based MPC with unique challenges when using the Fundamental Lemma for prediction, in particular inaccuracies due to noise or nonlinearities.
@@
    Transferring open-loop results to closed-loop formulations: Several effects in data-driven MPC are by now well-understood in open-loop optimal control formulations, but have not been researched extensively with respect to closed-loop guarantees, e.g., regularization. On the other hand, when applying data-driven MPC, typically the closed-loop performance is the main object of interest. Thus, transferring open-loop to closed-loop results is another promising future research avenue.